# Brill–Noether loci in genus $\leq 12$

### v1

@@ Abstract.
A refined Brill–Noether theory seeks to determine which linear series are admitted by a “general” curve in a particular Brill–Noether locus. However, as Brill–Noether loci are not irreducible in general, a coarse answer is given by the relative positions of Brill–Noether loci. Via an analysis of unstable Lazarsfeld–Mukai bundles on K3 surfaces, we distinguish Brill–Noether loci and provide expectations for the relative positions of Brill–Noether loci in general. Together with classical results, the refined Brill–Noether theory for curves of fixed gonality and on Hirzebruch surfaces, and explicit constructions, we identify the relative positions of all Brill–Noether loci in genus $g\leq 11$, and all but one locus in genus $12$.
@@ Introduction
when $\rho(g,r,d)<0$. A refined Brill–Noether theory has been recently established for curves of fixed gonality (addressing the $r=1$ case) by work of many authors in [7, 13, 35, 36, 39, 40, 55] (see also the excellent survey [34], and [18, 19, 42] for low gonality results), and for smooth curves on Hirzebruch surfaces by Larson–Vemulapalli in [41].
@@
The relative positions of Brill–Noether loci in genus $g\leq 6$ are given by trivial containments. For $7\leq g\leq 11$, the relative positions are identified in Figure 2 (genus $7$), Figure 3 (genus $8$), Figure 4 (genus $9$), Figure 5 (genus $10$), and Figure 7 (genus $11$).
Moreover, except for the potential containments $\mathcal{M}^{2}_{12,7}\stackrel{{\scriptstyle?}}{{\subset}}\mathcal{M}^{3}_{12,10}$ and $\mathcal{M}^{2}_{12,7}\stackrel{{\scriptstyle?}}{{\subset}}\mathcal{M}^{3}_{12,11}$, the relative positions of Brill–Noether loci in genus $12$ are identified in Figure 8. Additionally, we provide expectations for the relative positions of Brill–Noether loci in general via an analysis of unstable Lazarsfeld–Mukai bundles on K3 surfaces.
@@ Acknowledgments
The author would like to thank David Anderson, Asher Auel, Andrei Bud, Carl Lian, Montserrat Teixidor i Bigas, and Sameera Vemulapalli for helpful conversations, and particularly Isabel Vogt for discussions and insight, and Andreas Leopold Knutsen for helpful comments on an earlier draft of this paper. This research was partially conducted during the time that R.H. was supported by the National Science Foundation under Grant No. DMS-2231565.
@@
In Section 1, we recall background on Brill–Noether loci in Section 1.1, special linear series and classical results in Section 1.2, relevant results from the refined Brill–Noether theory for curves of fixed gonality in Section 1.3, Lazarsfeld–Mukai bundles and stability in Section 1.4, and curves on K3 surfaces in Section 1.5. In Section 2, we define admissible assignments for a terminal filtration of an unstable Lazarsfeld–Mukai bundle, give bounds on admissible assignments in Section 2.3, and summarize how they are used for distinguishing Brill–Noether loci in Section 2.4. In Section 3, we begin the proof of A, for genus 33336666 in Section 3.1, genus $7$ in Section 3.2, genus $8$ in Section 3.3, genus $9$ in Section 3.4, and genus $10$ in Section 3.5. In Section 4 we give some general containments and non-containments in higher genus, and return to the proof of A in genus $11$ in Section 4.2, and outline the relative positions in genus $12$ in Section 4.3. Finally, in Section 5, guided by curves on K3 surfaces and the results in low genus, we motivate and state questions regarding the reducibility and relative positions of Brill–Noether loci in general.
@@ 4.3. Genus 12
It becomes more difficult to distinguish all Brill–Noether loci in higher genus.
Using the results used in genus $\leq 11$, many containment and non-containments can be shown. Using Castelnuovo curves, a few additional relative positions can be pinned down, and the results are summarized in Figure 8.
${\mathcal{M}^{2}_{12,9}}$${\mathcal{M}^{3}_{12,11}}$${\mathcal{M}^{1}_{12,6}}$${\mathcal{M}^{2}_{12,8}}$${\mathcal{M}^{3}_{12,10}}$${\mathcal{M}^{1}_{12,5}}$${\mathcal{M}^{2}_{12,7}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,9}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,11}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{1}_{12,4}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{2}_{12,6}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,8}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,10}}$${\mathcal{M}^{1}_{12,3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{2}_{12,5}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,7}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,9}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{5}_{12,11}}$${\mathcal{M}^{1}_{12,2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Figure 8. Non-trivial containments of Brill–Noether loci in genus $12$. Dashed arrows signify unknown containment or non-containment.
The relative position of $\mathcal{M}^{2}_{12,7}$ is not known, as indicated by dashed arrows. We hope to establish the relative positions in genus $12$ in the future.
@@
This leaves only the potential containments of $\mathcal{M}^{2}_{12,7}$ into $\mathcal{M}^{3}_{12,10}$ or $\mathcal{M}^{3}_{12,11}$, for which constructions similar to 4.11 in genus $11$ may exist. An understanding the Brill–Noether theory of nodal plane curves with a small number of nodes may also address these containments.
@@
We note that the K3 surface techniques do not suffice to settle the potential containments on $\mathcal{M}^{2}_{12,7}$. Indeed, on a K3 surface with $\operatorname{Pic}(S)=\Lambda^{2}_{12,7}$, there are admissible assignments for Lazarsfeld–Mukai bundles of a $g^{3}_{10}$ and a $g^{3}_{11}$ on a smooth irreducible $C\in|H|$. There are two admissible assignments for a Lazarsfeld–Mukai bundle of a $g^{3}_{10}$ on $C\in|H|$, giving bundles
@@
Furthermore, given $D_{1}\in|H-L|$ and $D_{2}\in|L|$ be smooth curves, then $E_{D_{1},g^{1}_{4}}\oplus E_{L,g^{1}_{2}}$ is the Lazarsfeld–Mukai bundle of a $g^{3}_{11}$ on $C$. Perhaps $\mathcal{M}^{2}_{12,7}$ is in fact contained in $\mathcal{M}^{3}_{12,11}$, which is K3-expected.
@@ 4.4. Genus $\geq 13$
In general, one may provide a clear picture of the relative positions of Brill–Noether loci of not too large codimension For example, using the results of [6], in some range, $\operatorname{Pic}(S)=\Lambda^{r}_{g,d}$ has a unique flexible decomposition of $H$, and K3 surfaces suffice to identify non-containments. There may be a larger range still, where flexible decompositions are not unique, but an analysis of admissible assignments and other methods may be sufficient, which we pursue in forthcoming work. In particular, as in 3, it would be of interest to identify a range where there are no non-trivial containments among Brill–Noether loci, though this range would have to occur in large genus where $d_{max}(g,r)-2r>\left\lfloor\frac{g-1}{2}\right\rfloor$.
@@ 5.2. Gonality of curves in $\mathcal{M}^{r}_{g,d}$
In some range, it is perhaps expected that $K(g,r,d)=\min\{d-2r+2,~{}\left\lfloor(g+3)/2\right\rfloor\}$. However, already in genus $11$ and $12$ this does not hold. As observed in 4.6 and 4.14, $\mathcal{M}^{3}_{11,9}=\mathcal{M}^{2}_{11,6}\subset\mathcal{M}^{1}_{11,4}$, $\mathcal{M}^{3}_{12,9}\subset\mathcal{M}^{1}_{12,4}$, and $\mathcal{M}^{4}_{12,11}\subset\mathcal{M}^{1}_{12,4}$, and for higher genus one has 4.6 and 4.16.

### v2

@@ Abstract.
A refined Brill–Noether theory seeks to determine which linear series are admitted by a “general” curve in a particular Brill–Noether locus. However, as Brill–Noether loci are not irreducible in general, a coarse answer is given by the relative positions of Brill–Noether loci. Via an analysis of unstable Lazarsfeld–Mukai bundles on K3 surfaces, we distinguish Brill–Noether loci and provide expectations for the relative positions of Brill–Noether loci in general. Together with classical results, the refined Brill–Noether theory for curves of fixed gonality and on Hirzebruch surfaces, and explicit constructions, we identify the relative positions of all Brill–Noether loci in genus $g\leq 12$.
@@ Introduction
when $\rho(g,r,d)<0$. A refined Brill–Noether theory has been recently established for curves of fixed gonality (addressing the $r=1$ case) by work of many authors in [7, 13, 36, 39, 40, 55] (see also the excellent survey [34], and [18, 19, 42] for low gonality results), and for smooth curves on Hirzebruch surfaces by Larson–Vemulapalli in [41].
@@
The relative positions of Brill–Noether loci in genus $g\leq 6$ are given by trivial containments. For $7\leq g\leq 12$, the relative positions are identified in Figure 2 (genus $7$), Figure 3 (genus $8$), Figure 4 (genus $9$), Figure 5 (genus $10$), Figure 7 (genus $11$), and Figure 8 (genus $12$).
Additionally, we provide expectations for the relative positions of Brill–Noether loci in general via an analysis of unstable Lazarsfeld–Mukai bundles on K3 surfaces, and highlight additional techniques for distinguishing Brill–Noether loci, which may reduce the problem of identifying the relative positions to a small list where explicit constructions are required to prove containments.
@@ Acknowledgments
The author would like to thank David Anderson, Asher Auel, Andrei Bud, Hannah Larson, Carl Lian, Montserrat Teixidor i Bigas, and Sameera Vemulapalli for helpful conversations, and particularly to Dave Jensen and Isabel Vogt for discussions and insight, and Andreas Leopold Knutsen for helpful comments on an earlier draft of this paper. This research was partially conducted during the time that R.H. was supported by the National Science Foundation under Grant No. DMS-2231565.
@@
In Section 1, we recall background on Brill–Noether loci in Section 1.1, special linear series and classical results in Section 1.2, relevant results from the refined Brill–Noether theory for curves of fixed gonality in Section 1.3, Lazarsfeld–Mukai bundles and stability in Section 1.4, and curves on K3 surfaces in Section 1.5. In Section 2, we define admissible assignments for a terminal filtration of an unstable Lazarsfeld–Mukai bundle, give bounds on admissible assignments in Section 2.3, and summarize how they are used for distinguishing Brill–Noether loci in Section 2.4. In Section 3, we begin the proof of A, for genus 33336666 in Section 3.1, genus $7$ in Section 3.2, genus $8$ in Section 3.3, genus $9$ in Section 3.4, and genus $10$ in Section 3.5. In Section 4 we give some general containments and non-containments in higher genus, and return to the proof of A in genus $11$ and $12$ in Section 4.2 and Section 4.3. Finally, in Section 5, we motivate and state questions regarding the reducibility and relative positions of Brill–Noether loci in general.
@@ 4.3. Genus 12
It becomes more difficult to distinguish all Brill–Noether loci in higher genus. Using the results used in genus $\leq 11$, many containment and non-containments can be shown, and we fill in the reaming gaps with ad hoc constructions using Castelnuovo curves and linear systems of curves through nodal plane curves, as in genus $11$. The relative positions of Brill–Noether loci in genus $12$ are summarized in Figure 8.
${\mathcal{M}^{2}_{12,9}}$${\mathcal{M}^{3}_{12,11}}$${\mathcal{M}^{1}_{12,6}}$${\mathcal{M}^{2}_{12,8}}$${\mathcal{M}^{3}_{12,10}}$${\mathcal{M}^{1}_{12,5}}$${\mathcal{M}^{2}_{12,7}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,9}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,11}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{1}_{12,4}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{2}_{12,6}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,8}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,10}}$${\mathcal{M}^{1}_{12,3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{2}_{12,5}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{3}_{12,7}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{4}_{12,9}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}^{5}_{12,11}}$${\mathcal{M}^{1}_{12,2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
Figure 8. Non-trivial containments of Brill–Noether loci in genus $12$.
@@
This leaves only the potential containments of $\mathcal{M}^{2}_{12,7}$ into $\mathcal{M}^{3}_{12,10}$ or $\mathcal{M}^{3}_{12,11}$, for which we give a construction in 4.16 similar to 4.11. An understanding the Brill–Noether theory of nodal plane curves with a small number of nodes may also address similar containments. We first note the K3-expected containments of $\mathcal{M}^{2}_{12,7}$.
@@
We note that the containments $\mathcal{M}^{2}_{12,7}\subset\mathcal{M}^{3}_{12,10}\subset\mathcal{M}^{3}_{12,11}$ is K3-expected. Indeed, on a K3 surface with $\operatorname{Pic}(S)=\Lambda^{2}_{12,7}$, there are admissible assignments for Lazarsfeld–Mukai bundles of a $g^{3}_{10}$ and a $g^{3}_{11}$ on a smooth irreducible $C\in|H|$. There are two admissible assignments for a Lazarsfeld–Mukai bundle of a $g^{3}_{10}$ on $C\in|H|$, giving bundles
@@
Furthermore, given $D_{1}\in|H-L|$ and $D_{2}\in|L|$ be smooth curves, then $E_{D_{1},g^{1}_{4}}\oplus E_{L,g^{1}_{2}}$ is the Lazarsfeld–Mukai bundle of a $g^{3}_{11}$ on $C$.
We also note that the containment $\mathcal{M}^{2}_{12,7}\subset\mathcal{M}^{3}_{12,10}$ is also predicted by admissible fillings of tableaux; and also by considering more special curves $C\in\mathcal{M}^{2}_{12,7}$ whose image in $\mathbb{P}^{2}$ has a triple point which, after blowing up the triple point, lie on the Hirzebruch surface $\mathbb{F}_{1}$ and admit a $g^{3}_{10}$, as can be checked from [41, Theorem 1.5].
We thank Dave Jensen for the following containment.
Proposition 4.16.
There is a containment $\mathcal{M}^{2}_{12,7}\subset\mathcal{M}^{3}_{12,10}$.
Proof.
Let $C\in\mathcal{M}^{2}_{12,7}$. We first note that we may assume that the $g^{2}_{7}$ is basepoint free, as otherwise $C\in\mathcal{M}^{2}_{12,6}\subset\mathcal{M}^{3}_{12,9}\subset\mathcal{M}^{3}_{12,10}$ by 4.14. Thus, as $7$ is prime, we may assume that the $g^{2}_{7}$ maps $C$ birationally onto its image in $\mathbb{P}^{2}$, and that $C$ is general in the image of the Severi variety, and we may assume that the image of $C$ under the $g^{2}_{7}$ is a plane curve with $123=\binom{7-1}{2}-$ 12 nodes, which may be taken to be in general position. Taking the $\binom{4}{2}-1-1=4$ dimensional linear system of conics through $2$ of the nodes gives a $g^{3}_{10}$, as each conic intersects the image of $C$ in $10$ additional points. ∎
Theorem 4.17.
All non-trivial containments among Brill–Noether loci in genus $12$ are obtained via trivial containments and the arrows appearing in Figure 8.
@@ 4.4. Genus $\geq 13$
In general, one may provide a clear picture of the relative positions of Brill–Noether loci of not too large codimension, using some ad hoc constructions as in genus $11$ and $12$, where possible.
For example, using the results of [6], in some range, $\operatorname{Pic}(S)=\Lambda^{r}_{g,d}$ has a unique flexible decomposition of $H$, and K3 surfaces suffice to identify non-containments. There may be a larger range still, where flexible decompositions are not unique, but an analysis of admissible assignments and other methods may be sufficient, which we pursue in forthcoming work. In particular, as in 3, it would be of interest to identify a range where there are no non-trivial containments among Brill–Noether loci, though this range would have to occur in large genus where $d_{max}(g,r)-2r>\left\lfloor\frac{g-1}{2}\right\rfloor$.
@@ 5.2. Gonality of curves in $\mathcal{M}^{r}_{g,d}$
In some range, it is perhaps expected that $K(g,r,d)=\min\{d-2r+2,~{}\left\lfloor(g+3)/2\right\rfloor\}$. However, already in genus $11$ and $12$ this does not hold. As observed in 4.6 and 4.14, $\mathcal{M}^{3}_{11,9}=\mathcal{M}^{2}_{11,6}\subset\mathcal{M}^{1}_{11,4}$, $\mathcal{M}^{3}_{12,9}\subset\mathcal{M}^{1}_{12,4}$, and $\mathcal{M}^{4}_{12,11}\subset\mathcal{M}^{1}_{12,4}$, and for higher genus one has 4.6 and 4.18.